### v1

Abstract
A multiscale optimization framework for problems over a space of Lipschitz continuous functions is developed. The method solves a coarse-grid discretization followed by linear interpolation to warm-start project gradient descent on progressively finer grids. Greedy and lazy variants are analyzed and convergence guarantees are derived that show the multiscale approach achieves provably tighter error bounds at lower computational cost than single-scale optimization. The analysis extends to any base algorithm with iterate convergence at a fixed rate. Constraint modification techniques preserve feasibility across scales. Numerical experiments on probability density estimation problems, including geological data, demonstrate speedups of an order of magnitude or better.
keywords:
multiresolution analysis, continuous optimization, approximation theory
{MSCcodes}
65B99, 65D15, 90C59
1 Introduction
Optimization over function spaces arises in applications ranging from density estimation to signal reconstruction. We consider the general problem of optimizing an objective $\mathcal{L}:\mathcal{C}\to\mathbb{R}$ over continuous functions $f:\mathcal{D}\to\mathbb{R}$ defined on an $N$-dimensional domain $\mathcal{D}\subseteq\mathbb{R}^{N}$:
[…]
where $\mathcal{C}\subseteq\mathcal{F}:=\left\{f:\mathcal{D}\to\mathbb{R}\mid\text{$f$ is continuous}\right\}$ is a constraint set.
The space of all continuous functions $\mathcal{F}$ is too large to optimize directly. We therefore restrict $\mathcal{C}\subseteq\mathcal{F}$ to reduce the uncountably infinite dimensional problem in Eq.˜P to a countably infinite or finite problem amenable to computation. We focus on Lipschitz continuous solutions. This mild regularity condition holds automatically for continuous functions on compact domains and is needed to prove precise error bounds for our multiscale approach.
One common approach discretizes the problem over a finite grid:
[…]
where the tensor $X\in\mathbb{R}^{I_{1}\times\dots\times I_{N}}$ represents the function $f$ evaluated on a set of grid points. The tensor has $N$ modes, matching the dimension of $\mathcal{D}$, with $I_{n}$ grid points along the $n$th dimension. The discretized objective $\tilde{\mathcal{L}}$ and constraint set $\tilde{\mathcal{C}}$ correspond to their continuous counterparts.
Any discretization raises the question, How fine should be the discretization? How large should be the grid dimensions $I_{1},\dots,I_{N}$? When we have discretization choices to make, we favor finer grids for accuracy. When data arrives pre-discretized, we prefer to incorporate all available information. In either case, fine discretization (large $I_{n}$) is desirable. Fine discretizations, however, carry computational costs: most operations are slower and consume more memory as problem size grows, often superlinearly, such as with matrix-matrix multiplication [strassen_gaussian_1969]. We propose a multiresolution approach that achieves faster convergence on fine-scale problems. Our method is inspired by wavelets [benedetto_wavelets_1993] and multigrid [trottenberg_multigrid_2001] methods, and optimizes $f$ over progressively finer discretizations.
1.1 Related Work
Many approaches exist for optimizing over function spaces beyond simple discretizations. One class of methods restricts $\mathcal{C}$ to a Hilbert space with basis functions $\{f_{k}\}$ and solves for optimal weights $\{a_{k}\}$ in the representation $f(v)=\sum_{k}a_{k}f_{k}(v)$. Examples include Fourier features, wavelets [benedetto_wavelets_1993], and Legendre polynomials [trefethen_approximation_2019]. The calculus of variations provides an analytical framework for optimizing functionals, typically integrals involving the input function [gelfand_calculus_2000].
Finite parametrization offers another approach, exemplified by finding the mean $\mu$ and variance $\sigma^{2}$ in the family of normal distributions $f(v)=(2\pi\sigma)^{-1/2}\exp(-(v-\mu)^{2}/(2\sigma^{2}))$ [vershynin_HighDimensionalProbability_2018], or optimizing over weights $\{W_{k}\}$ and biases $\{b_{l}\}$ in fully connected neural networks [gurney_introduction_2018]. Non-parametric methods such as kernel density estimation [chen_tutorial_2017] and Gaussian mixture models [yan_learning_2023] provide yet another alternative. Finally, convex duality techniques transform infinite-dimensional problems into finite-dimensional counterparts [chuna_dual_2025].
We focus on uniformly spaced discretizations: the approach is simple, used across a variety of settings, easy to implement and analyze, and sufficient in practice and theory for Lipschitz functions. More sophisticated discretization schemes have been well studied in differential equations [stetter_analysis_1973], probability density estimation [chakraborty_generating_2015], machine learning [ramirez_gallego_data_2016], and approximation theory [trefethen_approximation_2019]. These schemes concentrate the grid where $f$ exhibits larger variations (for differentiable $f$, where the gradient magnitude is larger) or near domain endpoints. Many results in this paper extend to such alternative discretization schemes.
Our method synthesizes ideas from wavelets [benedetto_wavelets_1993], multigrid methods for partial differential equations [trottenberg_multigrid_2001], and multiresolution matrix factorization [kondor_multiresolution_2014]. At each scale $s$, we solve a distinct optimization problem ($\mathrm{P_{s}}$) that discretizes the continuous problem (P). Both the tensor dimension $X_{s}\in\mathbb{R}^{I_{s}^{1}\times\dots\times I_{s}^{N}}$ and the objective function $\tilde{\mathcal{L}}_{s}$ adapt to the scale: a coarse-scale objective involves fewer summands, different grid points, and modified constraints. This differs from approaches that merely subsample variables while keeping the objective structure fixed. By re-discretizing both the domain and the objective at each scale, our framework ensures that each subproblem $P_{s}$ properly approximates the continuous problem, not just its reduced fine-scale version.
Each resolution level provides a warm start for the next finer scale, typically with a smaller error than random initialization. Warm-start strategies appear throughout numerical optimization, from stochastic gradient descent [loshchilov_sgdr_2017] to first-order methods [adcock_restarts_2025].
Our coarse-scale optimization updates only a subset of coordinates (the free variables present at the current resolution) while finer-scale variables remain implicit. This connects to coordinate-wise optimization methods, including greedy descent [dhillon_nearest_2011], coordinate descent [nesterov_efficiency_2012], and block coordinate decent [xu_BlockCoordinateDescent_2013].
1.2 Contributions
We develop multiscale methods for optimizing over spaces of Lipschitz functions and establish the first theoretical framework for multiscale optimization in this setting. We propose two variants: a greedy approach that re-optimizes all variables at each scale, and a lazy approach that freezes coarse-scale values and optimizes only newly interpolated points. Both variants achieve faster convergence to tighter error bounds than gradient descent on a single fine discretization. The analysis extends to any optimization algorithm with iterate convergence at a fixed rate. We further develop constraint modification techniques that preserve feasibility across scales. Numerical tests illustrate the effectiveness of the multiscale approach on synthetic and realworld data for a distribution de-mixing problem.
1.3 Reproducible research
The data files and scripts used to generate the numerical results presented in this paper are available from the GitHub repository [Richardson_multiscale_paper_code].
2 The Multiscale Optimization Method
We introduce the multiscale method through a concrete example that demonstrates how it can exploit problem structure across discretization scales to reduce overall computation cost.
2.1 Motivating Example
Consider recovering a smooth probability density function from polynomial measurements. We formulate this as an inverse least-squares problem over differentiable densities $p$ on $\mathcal{D}=[-1,1]$:
[…]
where the constraints ensure $p$ is a valid probability density. The measurement operator $\mathcal{A}$ maps $p$ to an $M$-vector with elements
$\mathcal{A}(p)[m]=\left\langle a_{m},p\right\rangle\quad\text{for}\quad m=1,\dots,M,$
where the normalized Legendre polynomials are
$a_{m}(t)=\sqrt{\frac{2m+1}{2}}\sum_{k=0}^{m}\binom{m}{k}\binom{m+k}{k}\left(\frac{t-1}{2}\right)^{k}.$
The regularization term $\left\lVert p^{\prime}\right\rVert_{2}^{2}=\int_{-1}^{1}(p^{\prime}(t))^{2}\,dt$ encourages smoothness in the recovered density. We discretize this continuous problem on a uniform grid to obtain
[…]
where the vector $x_{1}=(x_{1}{[1]},x_{1}{[2]},\dots,x_{1}{[I_{1}]})$ represents the uniform discretization of $p$ on [1,1][-1,1] at the finest scale with grid spacing $\Delta t_{1}=2/(I_{1}-1)$:
$x_{1}{[i]}=p(t_{1}{[i]})\Delta t_{1},\qquad\text{with}\qquad t_{1}[i]=-1+2\frac{i-1}{I_{1}-1}.$
The discretized measurement operator $A_{1}$ and graph Laplacian $G_{1}$, which approximates the second derivative operator, are
$A_{1}[m,i]=a_{m}(t_{1}[i])\quad\text{and}\quad G_{1}=\frac{1}{\Delta t_{1}^{3}}\begin{bmatrix}1&-1&&&\\ -1&2&-1&&\\ &\ddots&\ddots&\ddots&\\ &&-1&2&-1\\ &&&-1&1\end{bmatrix}$.
We return to this example in Section˜2.2, after introducing the general multiscale framework, to demonstrate how the method applies to this specific problem.
Figure 1: Comparison of multiscale vs single-scale approach for the motivating example from Section˜2.1. Dots represent the median total time in milliseconds over $100$ trials; shaded regions represent the $5$th and $95$th percentile times. The multiscale approach improves algorithm speed as problem size grows. Details are provided in Section˜2.3.
2.2 Method Overview
We solve the continuous optimization problem Eq.˜P via progressive grid refinement. Starting from a coarse discretization, we solve the discretized problem at each scale, interpolate the solution to the next finer grid, and use it to warm-start the optimization. Algorithm˜1 formalizes this refinement process.
Algorithm 1 High-level multiscale method
[…]
1:  Solve $P_{S}$ with initial point $X_{S}^{0}$ to obtain $X_{S}$.
2:for scales $s=S-1,S-2,\dots,1$ do
3:   Compute $\underline{X}_{s+1}$ by interpolating $X_{s+1}$.
4:   Solve $P_{s}$ with initial point $X_{s}^{0}=\underline{X}_{s+1}$ to obtain $X_{s}$.
5:end for
6:  Construct piecewise linear approximation $\hat{f}_{X_{1}}$ from $X_{1}$.
7:return $\hat{f}_{X_{1}}$, an approximate solution to $P$.
We index discretization levels by a scale parameter $s$, where $s=1$ denotes the finest discretization and increasing $s$ corresponds to progressively coarser grids. At scale $s$, the discretized problem becomes
[…]
where $X_{s}$ is a real-valued $N$-mode tensor with dimension $(I_{s}^{1}\times\dots\times I_{s}^{N})$, and $\tilde{\mathcal{L}}_{s}$ and $\tilde{\mathcal{C}}_{s}$ denote the objective and constraint set discretized at scale $s$.
We return to the motivating example from Section˜2.1 to demonstrate how the general framework applies. Recall that we recover a probability density from Legendre polynomial measurements, discretized at the finest scale as Eq.˜2. At scale $s$, the discretized problem becomes
[…]
where $x_{s}$ is an $I_{s}=2^{s-1}+1$ vector that discretizes $p$ at scale $s$ with grid spacing $\Delta t_{s}=2^{s-1}\Delta t_{1}$. The constraint $\lVert x_{s}\rVert_{1}=2^{s-1}$ scales with the grid spacing to preserve the continuous constraint $\int p(t)\,dt=1$ across all scales (see Section˜3). The scaled measurement operator $A_{s}$ and graph Laplacian $G_{s}$ are defined as
$A_{s}=2^{s-1}A_{1}[:,1:2^{s-1}:I_{1}]\quad\text{and}\quad G_{s}=2^{1-s}G_{1}[1:2^{s-1}:I_{1},1:2^{s-1}:I_{1}],$
where the slicing notation $1:2^{s-1}:I_{1}$ selects every $(2^{s-1})$-th point from the finest grid. The prefactors $2^{s-1}$ and $2^{1-s}$ ensure the two main quantities in the loss remain balanced at each scale:
$A_{s}x_{s}\approx A_{1}x_{1}\quad\text{and}\quad x_{s}^{\top}G_{s}x_{s}\approx x_{1}^{\top}G_{1}x_{1}.$
This balancing ensures each scaled problem is regularized similarly by $\lambda$.
2.3 Motivating Example: Numerics
To emphasise the concepts behind multiscale, we compare an instance of the multiscale method (see Algorithm˜2) against single scale optimization on the motivating example from Section˜2.1. Figure˜1 illustrates the improvement in computation time versus problem size achieved by the multiscale approach. Additional numerics and details are provided in Section˜7.
We compare total runtime for $S=3,4,\dots,12$, which yields finest discretizations with $I_{1}=2^{S}+1$ points. We initialize by uniformly discretizing the polynomial $p(t)=-2.625t^{4}-1.35t^{3}+2.4t^{2}+1.35t+0.225$ on [1,1][-1,1] to obtain the initial density approximation $x_{S}[i]=f(t_{S}[i])\Delta t_{S}$ at the coarsest scale.
At each scale $s>1$, we perform one ($K_{s}=1$) iteration; at the finest scale ($s=1$), and we iterate until the objective is within $5\%$ of its optimal value. We generate Legendre measurements for $m=1,\dots,5$ using the measurement operator $A_{1}$ from Section˜2.1 and add $5\%$ Gaussian noise to obtain $y$. The regularization parameter $\lambda=10^{-4}$ balances data fidelity against smoothness.
Figure 2: Typical progression of iterates $x_{s}^{k}$ using single scale (top row) and multiscale (bottom row) approachs for the motivating example from Section˜2.1. The clean ground truth is plotted behind in the thick dashed lined.
In Fig.˜2, we show the typical convergence of iterates $x_{s}^{k}$ for single scale and multiscale approaches. We run the motivating example from Section˜2.1 with $2^{10}+1$ points at the finest scale $s=1$. The top-left plot shows every 25th iterate $k$ starting with the initialization at $k=0$ and ending at $k=250$ when running projected gradient descent at the finest scale ($s=1$). The top-right plot shows every 250th iterate $k$, ending with the final iterate $k=4988$ which satisfies the stopping criteria $\tilde{\mathcal{L}}_{1}(x_{1}^{k})<0.00024$. The bottom-left plot shows $x_{s}^{k}$ at each scale from the coarsest scale $s=10$ to the second finest $s=2$, and the iterate at the finest scale $s=1$ after $k=25$ iterations. We can make a fair comparison between $x_{1}^{25}$ on the top-left and bottom-left plots since they are both the result of running $k=25$ iterations of projected gradient descent at the finest scale. Using multiscale results in a smoother iterate since $x_{1}^{0}$ was constructed from the chain of interpolations, rather than the completely random initialization used by the single scale approach. Because of this, the bottom-left plot highlights multiscale only needs $k=387$ iterations at the finest scale to also achieve the same stopping criteria $\tilde{\mathcal{L}}_{1}(x_{1}^{k})<0.00024$.
We call multiscale’s tendency to smooth iterates an implicit regularization. This helps smooth solutions by carrying information from farther away points to speed up convergence. In this problem, the graph laplacian regularizer only looks at the immediate neighbouring points. So working at coarser scales lets us compare points that are not immediately neighbouring at the finest scale.
Figure 3: Typical loss convergence using single scale (dashed blue) and multiscale (solid orange) approachs for the motivating example from Section˜2.1. The left plot shows the first millisecond, and the right plot shows the first 20 milliseconds. The left plot also highlights in grey regions that are not from projected gradient descent steps. These are the result of additional overhead with the multiscale method such as interpolation, allocations, and extra function calls.
In Fig.˜3, we show the typical loss convergence for the single scale and multiscale methods. These are shown as a function of time rather than number of iterations. Comparing iterations at different scales is not fair since we expect one projected gradient step to be cheaper at coarser scales. We also want to highlight the minimal effect of interpolation, allocations, and additional function calls. The loss $\tilde{\mathcal{L}}_{s}(x_{s}^{k})$ for the multiscale method is calculated at the respective scale $s$ of the iterate, whereas the single scale only calculates the the loss $\tilde{\mathcal{L}}_{1}(x_{1}^{k})$ at the finest scale $s=1$. This explains why the loss fluctuates up and down during interpolations for coarser scales $s$ when the approximation $\tilde{\mathcal{L}}_{s}$ for $\mathcal{L}$ is worse. The coarsest scales also take longer than expected because of the additional function call overhead. Despite these drawbacks, multiscale more than makes up for these by reducing the loss significantly faster once iterating on the finer scales. This is make clear by the right plot in Fig.˜3.
It could be argued that for optimal convergence rates with a first order method, we should not be using a fixed stepsize based on the global Lipschitz constant [nesterov-smooth-2018]. Using accelerated methods would almost certainly accelerate convergence and not require almost $5000$ iterations. But any sort of acceleration could also be used to speed up convergence for the multiscale method which already has a leg-up since the fine scale iterations start closer to the solution.
2.4 Descritization, Coarsening, and Interpolation
Consider a nested sequence of discretizations
(4) $\mathcal{D}_{S}\subset\mathcal{D}_{S-1}\subset\dots\subset\mathcal{D}_{1}\subset\mathcal{D}\subseteq\mathbb{R}^{N}$}
of the continuous domain $\mathcal{D}$. Scale $s=S$ represents the coarsest grid. Each discretization $\mathcal{D}$ contains $\left\lvert\mathcal{D}_{s}\right\rvert=\prod_{n=1}^{N}I_{s}^{n}$ points. The tensor $T$ with dimension $(I_{s}^{1}\times\dots\times I_{s}^{N})$ contains the discretization points from $\mathcal{D}_{s}$. The optimization variable $X_{s}$ represents the function values at these points:
$X_{s}[i_{1},\dots,i_{N}]=f(T_{s}[i_{1},\dots,i_{N}]).$
Unless otherwise stated, we restrict the remainder of the paper to scalar functions $f:[\ell,u]\to\mathbb{R}$ on the 1-dimensional domain $\mathcal{D}=[\ell,u]$ with uniform discretization
$x_{s}[i]=f(t_{s}[i]),\qquad\text{with}\qquad t_{s}[i]=\ell+\frac{i-1}{I_{s}-1}(u-\ell).$
This specialization simplifies our exposition and is without loss of generality because the multiscale framework extends to tensor-valued functions by vectorizing $X_{s}\in\mathbb{R}^{I_{1}\times\dots\times I_{N}}$ as $x_{s}\in\mathbb{R}^{I_{1}\cdots I_{N}}$ and adapting interpolation operators accordingly. Section˜7 demonstrates this extension.
[…]
Definition 2.1 (Dyadic Coarsening).
Dyadic coarsening maps a vector $x\in\mathbb{R}^{I}$ to $\overline{x}\in\mathbb{R}^{\lfloor(I+1)/2\rfloor}$ by selecting odd-indexed entries:
$\overline{x}[i]=x[2i-1],\qquad i=1,\ldots,\lfloor(I+1)/2\rfloor.$
Definition 2.2 (Midpoint Linear Interpolation).
Midpoint linear interpolation inserts a linearly interpolated point between each adjacent pair of entries in $x\in\mathbb{R}^{I}$ to produce $\underline{x}\in\mathbb{R}^{2I-1}$:
$\underline{x}[i]=\begin{cases}x[\frac{i+1}{2}]&\text{if $i$ is odd,}\\ \frac{1}{2}\left(x[\frac{i}{2}]+x[\frac{i}{2}+1]\right)&\text{if $i$ is even.}\end{cases}$
The operators in the coarsening and interpolation definitions translate between scales $s$ with the notation
$\overline{x}_{s}=x_{s+1}\qquad\text{and}\qquad\underline{x}_{s}=x_{s-1}.$
The overline operator $\overline{(\cdot)}$ maps $x_{s}$ to the coarser scale $s+1$; the underline operator $\underline{(\cdot)}$ maps to the finer scale s1s-1; see Fig.˜4.
When analyzing the multiscale algorithm, we need to refer to the newly added interpolated points. We call this vector the free variables $x_{(s)}$, a term justified in Section˜2.5.
Definition 2.3 (Vector of Free Variables).
Given the $(I_{s+1})$-vector $x_{s+1}$ and its interpolation $\underline{x}_{s+1}=x_{s}$, the $(I_{s+1}-1)$-vector of free variables has entries
$x_{(s)}[i]=\frac{1}{2}(x_{s+1}[i]+x_{s+1}[i+1]),\quad i=1,\dots,I_{s+1}-1.$
These are the newly added points in $x_{s}$ when interpolating $x_{s+1}$. At the coarsest scale $s=S$, we set $x_{(S)}=x_{S}$.
Figure 4: Example uniform discretization of an interval with $S=3$ scales. Each scale $s$ has $2^{S-s+1}+1$ points. The free variables $x_{(s)}$ (magenta squares) resemble a “W-cycle” from multigrid [trottenberg_multigrid_2001].
[…]
Definition 2.4 (Piecewise Linear Approximation).
Let the $I$-vector $x$ represent (possibly approximate) samples of a function $f$ at locations $t_{1}<\dots<t_{I}$. The piecewise linear approximation constructed from $x$ is
$\hat{f}_{x}(t)=\begin{cases}\frac{{x[2]-x[1]}}{t_{2}-t_{1}}(t-t_{1})+x[1]&\text{if $t_{1}\leq t\leq t_{2}$,}\\ \qquad\qquad\vdots\\ \frac{{x[I]-x[I-1]}}{t_{I}-t_{I-1}}(t-t_{I-1})+x[I-1]&\text{if $t_{I-1}\leq t\leq t_{I}$.}\end{cases}$
Remark 2.5.
The subscript in $\hat{f}_{x}$ identifies the vector $x$ that defines the approximation. When $x$ contains samples from a function, that is, $x[i]=f(t_{i})$, we omit the subscript and write $\hat{f}(t)$.
2.5 Greedy and Lazy MultiScale Algorithms
We present two variants of the multiscale method outlined in Section˜2.2 for solving Eq.˜P: the greedy variant (Algorithm˜2), which optimizes all grid points at each scale, and the lazy variant (Algorithm˜3), which optimizes only the newly interpolated points at each finer scale. Both variants apply an update rule $x^{k+1}=U(x^{k})$ at each scale, and our convergence analysis holds for any such update rule that achieves q-linear iterate convergence.
Figure 5: One hypothetical round of multiscale optimization. In all four subplots, the solid blue curve represents the solution to the continuous problem. The orange dots represent a) a random coarse initialization with four points, b) the solution at the coarse scale, c) the interpolation of the coarse solution to the fine grid, and d) the solution at the fine scale.
Definition 2.6 (Q-Linear Iterate Convergence).
An iterative algorithm with update rule $x^{k+1}=U(x^{k})$ has global q-linear iterate convergence with rate $q\in[0,1)$ when
$\left\lVert x^{k}-x^{*}\right\rVert_{2}\leq(q)^{k}\left\lVert x^{0}-x^{*}\right\rVert_{2}$
for any feasible initialization $x^{0}\in\tilde{\mathcal{C}}$ and some minimizer $x^{*}$ of $\tilde{\mathcal{L}}(x)$ over $\tilde{\mathcal{C}}$ that may depend on $x^{0}$.
Figure˜5 illustrates one round of multiscale optimization. Suppose the solution to the continuous optimization problem is
$f(t)=-84t^{4}+146.4t^{3}-74.4t^{2}+12t$
on the interval $0\leq t\leq 1$. The figure displays the multiscale flow: initialize at the coarse scale with five points, optimize at the coarse scale, interpolate to the fine grid (adding four points), and optimize at the fine scale. In this particular example, the greedy and lazy variants share this structure but differ in the final optimization: greedy optimizes all nine fine-scale points, while lazy optimizes only the four newly interpolated points.
Our software implementation, described in Section˜7, uses projected gradient descent as the update rule, which achieves $q$-linear iterate convergence for smooth and strongly convex objectives $\tilde{\mathcal{L}}(x)$ with closed and convex constraints $\tilde{\mathcal{C}}$ [beck_first-order_2017, Sec. 10.2]. This update rule takes the form
(5) $x^{k+1}\leftarrow U(x^{k})=P_{\tilde{\mathcal{C}}}\left(x^{k}-\alpha\nabla\tilde{\mathcal{L}}(x^{k})\right)$
where
[…]
denotes the projection of $x$ onto $\tilde{\mathcal{C}}$. The standard analysis of projected gradient descent gives the following $q$-linear rate for strongly-convex objectives.
Lemma 2.7 (Descent Lemma [nesterov-smooth-2018, Theorem 2.1.15]).
Projected gradient descent with stepsize $\alpha$ has iterate convergence with rate
$q(\alpha)=\sqrt{1-\frac{2\alpha\mathcal{S}_{\tilde{\mathcal{L}}}\mu_{\tilde{\mathcal{L}}}}{\mathcal{S}_{\tilde{\mathcal{L}}}+\mu_{\tilde{\mathcal{L}}}}}$.
[…]
$q(\alpha)=\frac{c-1}{c+1}$
for a condition number of $c=\mathcal{S}_{\tilde{\mathcal{L}}}/\mu_{\tilde{\mathcal{L}}}$.
Although we can prescribe the number of iterations $K_{s}$ at each scale in advance, in practice we may also use additional convergence criteria (such as a small gradient) at each scale in addition to this predetermined maximum.
Algorithm 2 Greedy multiscale method
1:  Randomly initialize $x_{s}^{0}$ at the coarsest scale $s=S$.
[…]
3:for scales $s=S-1,S-2,\dots,1$ do
4:   Interpolate $x_{s+1}^{K_{s+1}}$ to obtain $\underline{x}_{s+1}^{K_{s+1}}$ according to Definition˜$2.$
5:   Initialize the next finest scale $x_{s}^{0}=\underline{x}_{s+1}^{K_{s+1}}$ using the previous scale’s interpolated solution.
[…]
7:end for
[…]
The lazy variant freezes previously computed entries of $x$ and adjusts only the newly interpolated values $x_{(s)}$ (Definition˜2.3) during the algorithm update step. Figure˜4 illustrates these free variables. Algorithm˜3 implements this approach, differing from Algorithm˜2 only in lines 1, 2, and 6.
Algorithm 3 Lazy multiscale method
1:  Randomly initialize $x_{(s)}^{0}$ (Definition˜2.3) at the coarsest scale $s=S$.
[…]
3:for scales $s=S-1,S-2,\dots,1$ do
4:   Interpolate $x_{s+1}^{K_{s+1}}$ to obtain $\underline{x}_{s+1}^{K_{s+1}}$ according to Definition˜$2.$
5:   Initialize the next finest scale $x_{s}^{0}=\underline{x}_{s+1}^{K_{s+1}}$ using the previous scale’s interpolated solution.
[…]
7:end for
[…]
3 Discretizing Constraints
The multiscale approach requires discretizing Eq.˜P at multiple scales. Given a discretization of the continuous constraint $\mathcal{C}$ from (P) to the finest scale $\tilde{\mathcal{C}}_{1}$ in ($\mathrm{P_{s}}$), we must determine the appropriate discretized constraint sequence $\tilde{\mathcal{C}}_{s}$ at coarser scales $s=S-1,S-2,\dots,1$.
Pointwise constraints $\mathcal{C}=\left\{f:\mathcal{D}\to\mathbb{R}\mid f(t)\in\mathcal{X}\right\}$ transfer verbatim to any scale: $\tilde{\mathcal{C}}_{s}=\left\{x_{s}\in\mathbb{R}^{I_{s}}\mid x[i]\in\mathcal{X}\right\}.$ However, linear and moment constraints require careful scaling to maintain consistency across discretization levels.
Consider a $p$-norm constraint on $f$:
[…]
The natural discretization
[…]
correctly approximates $\mathcal{C}$ when we scale the entries of $x_{1}$ by the grid size:
$x_{1}[i]=f(t_{1}[i])(\Delta t)^{1/p}.$
With this scaling, the discrete norm approximates the continuous norm:
$c^{p}\coloneqq\left\lVert x_{1}\right\rVert_{p}^{p}=\sum_{i=1}^{I_{1}}f(t_{1}[i])^{p}\Delta t\approx\int_{\mathcal{D}}f(t)^{p}\,dt=\left\lVert f\right\rVert_{p}^{p},$
as $I_{1}\to\infty$.
At coarse scale $s$, we modify the constraint to
[…]
This scaling preserves the constraint across scales. The Lipschitz property of $f$ ensures neighboring entries in $x_{s}$ remain similar, so
$\frac{I_{1}}{I_{s}}\sum_{i=1}^{I_{s}}(x_{s}[i])^{p}\approx\sum_{i=1}^{I_{1}}(x_{1}[i])^{p},$
which gives
$c^{p}=\frac{I_{1}}{I_{s}}\left\lVert x_{s}\right\rVert_{p}^{p}\approx\left\lVert x_{1}\right\rVert_{p}^{p}.$
Linear constraints require analogous scaling. For $L_{g_{k}}$-Lipschitz functions $g_{k}:\mathcal{D}\to\mathbb{R}$, the continuous constraint
[…]
discretizes at scale $s$ as
[…]
where
$A_{s}[k,i]=g_{k}(t_{s}[i])(\Delta t)^{1/2}\quad\text{and}\quad x_{s}[i]=f(t_{s}[i])(\Delta t)^{1/2}.$
We bound the discretization error to justify our constraint modification across scales. We assume grid points double at each scale refinement ($I_{s}/I_{s+1}=2$) and use a uniform grid $\{t_{i}\}$ on the domain $\mathcal{D}=[\ell,u]$ to sample an $L_{f}$-Lipschitz function $f$ according to $x[i]=f(t_{i})$. The following proposition quantifies how linear constraints transform under coarsening.
Proposition 3.1 (Single Linear Constraint Scaling).
Let $a\in\mathbb{R}^{I}$ discretize a bounded $L_{g}$-Lipschitz function $g:[\ell,u]\to\mathbb{R}$, and let $x$ discretize a bounded $L_{f}$-Lipschitz function $f$. Define the Lipschitz constant for the product function $fg$ as $L_{fg}=L_{f}\left\lVert g\right\rVert_{\infty}+L_{g}\left\lVert f\right\rVert_{\infty}$.
If $\langle a,x\rangle=b$, then the coarsened constraint satisfies the bounds
$\left\lvert\left\langle\overline{a},\overline{x}\right\rangle-\tfrac{1}{2}b\right\rvert$ $\leq\tfrac{1}{4}L_{fg}(u-l)\left(\frac{I}{I-1}\right)$ $\quad\text{for even $I$},$
$\left\lvert\left\langle\overline{a},\overline{x}\right\rangle-\tfrac{1}{2}\frac{I+1}{I}b\right\rvert$ $\leq\tfrac{1}{2}L_{fg}(u-l)$ $\quad\text{for odd $I$}.$
Proof 3.2.
See LABEL:sec-linear-constraint-scaling-proof.
Proposition˜3.1 extends naturally to multiple linear constraints, as described in the following corollary.
Corollary 3.3 (Matrix Linear Constraint Scaling).
Let $g_{k}:[\ell,u]\to\mathbb{R}$ be bounded $L_{g_{k}}$-Lipschitz functions for $k\in[K]$. Let $A\in\mathbb{R}^{K\times I}$ be the discretization $A[k,i]=g_{k}(t[i])$, and define $\overline{A}=A[:,\texttt{begin}\mathbin{:}2\mathbin{:}\texttt{end}]$ as the $(K\times\left\lceil I\right\rceil_{e}/2)$-submatrix with every other column removed, where $\left\lceil I\right\rceil_{e}$ rounds up to the nearest even integer. Define $L_{fg}=\max_{k\in[K]}(L_{fg_{k}})=\max_{k\in[K]}(L_{f}\left\lVert g_{k}\right\rVert_{\infty}+L_{g_{k}}\left\lVert f\right\rVert_{\infty})$ as the maximum Lipschitz constant of the product functions $\{fg_{k}\}_{k\in[K]}$.
If $Ax=b\in\mathbb{R}^{K}$, then
$\left\lVert\overline{A}\overline{x}-\tfrac{1}{2}b\right\rVert_{2}$ $\leq\sqrt{K}\cdot\tfrac{1}{4}L_{fg}(u-l)\left(\frac{I}{I-1}\right)$ $\quad\text{for even $I$},$
$\left\lVert\overline{A}\overline{x}-\frac{I+1}{I}\tfrac{1}{2}b\right\rVert_{2}$ $\leq\sqrt{K}\cdot\tfrac{1}{2}L_{fg}(u-l)$ $\quad\text{for odd $I$}.$
Proof 3.4.
See LABEL:sec-matrix-linear-constraint-proof.
The 1-norm constraint admits analogous scaling bounds.
Proposition 3.5 (L1-norm Constraint Scaling).
Let $x\in\mathbb{R}^{I}$ discretize a bounded $L_{f}$-Lipschitz function $f$. If $\lVert x\rVert_{1}=b$, then the coarsened constraint satisfies the bounds
$\left\lvert\left\lVert\overline{x}\right\rVert_{1}-\tfrac{1}{2}b\right\rvert$ $\leq\tfrac{1}{4}L_{f}(u-l)\left(\frac{I}{I-1}\right)$ $\quad\text{for even $I$},$
$\left\lvert\left\lVert\overline{x}\right\rVert_{1}-\frac{I+1}{I}\tfrac{1}{2}b\right\rvert$ $\leq\tfrac{1}{2}L_{f}(u-l)$ $\quad\text{for odd $I$}.$
Proof 3.6.
The proof is similar to the proof of Proposition˜3.1 (see LABEL:sec-linear-constraint-scaling-proof) by considering the vector $a=\mathrm{sign}(x)$, and using the Lipschitz constant $L_{f}$ for $f$ to bound $\left\lvert x[i]-x[i+1]\right\rvert$ in place of $L_{fg}$.
Proposition 3.5 specializes to a clean asymptotic behavior in the normalized case. For $z=x/\lVert x\rVert_{1}$, the bound for even $I$ becomes
$\left\lvert\left\lVert\overline{z}\right\rVert_{1}-\tfrac{1}{2}\right\rvert\leq\tfrac{1}{4b}L_{f}(u-l)\left(\frac{I}{I-1}\right).$
As the discretization refines ($I\to\infty$), we have $b=\lVert x\rVert_{1}\to\infty$, so the bound vanishes and $\left\lVert\overline{z}\right\rVert_{1}\to 1/2$. The same limit holds for odd $I$, and also extends to other $p$-norms.
4 Convergence of a Multiscale Method
Both the standard and multiscale approaches execute many iterations at the finest scale and converge because, by assumption, they both apply a convergent algorithm at the finest scale. The multiscale achieves faster convergence by warm starting the fine-scale iterations with solutions from coarser scales. The next subsections develop this analysis.
Section˜4.1 establishes tight bounds on the error between smooth and strongly convex Lipschitz functions and their linear interpolation. Section˜4.2 proves convergence when we solve the problem along the entire new discretization at each scale. Section˜4.3 proves convergence when we freeze previously computed points and solve only at the newly interpolated points for each scale.
4.1 Functional Analysis Lemmas
[…]
The following lemma provides the foundation for analyzing linear interpolations of Lipschitz functions and, ultimately, our multiscale algorithm.
Lemma 4.1 (Lipschitz Function Interpolation).
Let $f:\mathbb{R}^{N}\to\mathbb{R}$ be $L_{f}$-Lipschitz. For all $a,b\in\mathbb{R}^{N}$ and $\lambda_{1},\lambda_{2}\geq 0$, $\lambda_{1}+\lambda_{2}=1$, the error between $f$ evaluated at a convex combination of $a$ and $b$ and the linear interpolation of $f(a)$ and $f(b)$ satisfies
(6) $\left\lvert f(\lambda_{1}a+\lambda_{2}b)-(\lambda_{1}f(a)+\lambda_{2}f(b))\right\rvert\leq 2L_{f}\lambda_{1}\lambda_{2}\left\lVert a-b\right\rVert_{2}.$}
Proof 4.2.
See LABEL:sec-lipschitz-interpolation-proof.
Figure 6: Example for Lemma˜4.1 with the function $x=f(t)=t-\cos(3\pi t)$ on the interval $\ell,u]=[0,1]$. The function $f$ must lie within the dashed parallelogram since it is $L_{f}=1+3\pi$ Lipschitz. Lemma˜4.1 uses the parallelogram constraint to bound the distance between $f$ and the linear interpolation $\lambda_{1}f(a)+\lambda_{2}f(b)$. Lemma˜4.3 shows that the bound in Eq.˜6 is tight for any given $\lambda_{1},\lambda_{2}\geq 0$ with $\lambda_{1}+\lambda_{2}=1$.
Since Lemma˜4.1 is central to the analysis of linear interpolations of Lipschitz functions and ultimately our multiscale algorithm, we require a tight bound on the interpolation error. The bound in (6) is optimal among all bounds independent of the endpoint values $f(a)$ and $f(b)$, as demonstrated by the explicit construction in the proof of Lemma˜4.3.
Lemma 4.3 (Lipschitz Interpolation Tightness).
Given $\lambda_{1},\lambda_{2}\geq 0$ with $\lambda_{1}+\lambda_{2}=1$, an interval $\ell,u]$, and constant $L_{f}>0$, there exists an $L_{f}$-Lipschitz function on $\ell,u]$ such that Eq.˜6 holds with equality.
Proof 4.4.
Construct
$f(t)=L_{f}\left\lvert t-(\lambda_{1}\ell+\lambda_{2}u)\right\rvert$.
This function is $L_{f}$-Lipschitz with end-point values $f(\ell)=L_{f}\lambda_{2}\left\lvert\ell-u\right\rvert$ and $f(u)=L_{f}\lambda_{1}\left\lvert\ell-u\right\rvert$, and $f(\lambda_{1}\ell+\lambda_{2}u)=0$. Therefore,
$\left\lvert\lambda_{1}f(\ell)+\lambda_{2}f(u)-f(\lambda_{1}\ell+\lambda_{2}u)\right\rvert$
$\quad\qquad=\big\lvert\lambda_{1}L_{f}\lambda_{2}\left\lvert\ell-u\right\rvert+\lambda_{2}L_{f}\lambda_{1}\left\lvert\ell-u\right\rvert\big\rvert$
$\quad\qquad=2L_{f}\lambda_{1}\lambda_{2}\left\lvert\ell-u\right\rvert$.
To bound the error introduced when interpolating a coarse-scale solution to a finer scale, we apply Lemma˜4.1 repeatedly with centre point interpolation ($\lambda_{1}=\lambda_{2}=1/2$). This bounds the error between an exact fine-scale discretization $x_{s}^{*}[j]=f(t_{s}[j])$ and the linear interpolation $x_{s}=\underline{x}_{s+1}$ obtained from a coarser discretization $x_{s+1}[i]=f(t_{s+1}[i])$. See Fig.˜7 (top).
Lemma 4.5 (Exact Interpolation).
Let $f:[\ell,u]\to\mathbb{R}$ be $L_{f}$-Lipschitz continuous. On the interval $\ell,u]$, let $t_{s}$ be a fine grid with J=2I1J=2I-1 points, and $t_{s+1}$ be a coarse grid with $I$ points. Discretize the function exactly at both scales to obtain vectors $x^{*}_{s}\in\mathbb{R}^{J}$ and $x^{*}_{s+1}\in\mathbb{R}^{I}$ where $x^{*}_{s}[j]=f(t_{s}[j])$ and $x^{*}_{s+1}[i]=f(t_{s+1}[i])$. Let $x_{s+1}=x^{*}_{s+1}\in\mathbb{R}^{I}$ and linearly interpolate to obtain $x_{s}=\underline{x}_{s+1}\in\mathbb{R}^{J}$ where
$\underline{x}_{s+1}[j]=\begin{cases}x_{s+1}\left[\frac{j+1}{2}\right]&\text{if $j$ is odd,}\\ \frac{1}{2}\left(x_{s+1}\left[\frac{j}{2}\right]+x_{s+1}\left[\frac{j}{2}+1\right]\right)&\text{if $j$ is even.}\end{cases}$
Then the difference between the interpolated $x_{s}$ and exact values $x_{s}^{*}$ is bounded by
(7) $\left\lVert x_{s}-x_{s}^{*}\right\rVert_{2}\leq L_{f}\frac{1}{2\sqrt{I-1}}\left\lvert u-\ell\right\rvert$.}
Proof 4.6.
See LABEL:sec-exact-interpolation-proof.
When the coarse discretization contains error, the interpolation analysis extends naturally. Suppose we interpolate not from exact values $x_{s+1}[i]=x_{s+1}^{*}[i]=f(t_{s+1}[i])$, but from approximate values $x_{s+1}[i]=x_{s+1}^{*}[i]+\delta[i]=f(t_{s+1}[i])+\delta[i]$, where $\delta$ is an $I$-vector that represents the error at the coarse scale. Fig.˜7 (bottom) visualizes this scenario.
Lemma 4.7 (Inexact Interpolation).
Consider the setup of Lemma˜4.5 with one modification: the coarse discretization $x_{s+1}=x^{*}_{s+1}+\delta$ contains error $\delta\in\mathbb{R}^{I}$ before we interpolate to obtain $x_{s}=\underline{x}_{s+1}$. The difference between the interpolated values $x_{s}$ and exact values $x_{s}^{*}$ satisfies
(8) $\left\lVert x_{s}-x_{s}^{*}\right\rVert_{2}\leq\sqrt{2}\left\lVert\delta\right\rVert_{2}+L_{f}\frac{1}{2\sqrt{I-1}}\left\lvert u-\ell\right\rvert$.}
Proof 4.8.
See LABEL:sec-inexact-interpolation-proof.
This bound decomposes into two terms corresponding to distinct error sources. The first term captures the error from interpolating approximate values $x_{s+1}=x^{*}_{s+1}+\delta$ rather than exact values $x^{*}_{s+1}$. The second term captures the linear interpolation error as in Lemma˜4.5.
Figure 7: Illustration of Lemma˜4.5 (top) and Lemma˜4.7 (bottom) for the function $f(t)=72t^{3}-80t^{2}+10t+1$. Both lemmas bound the error between the true fine-scale discretization $x^{*}_{s}$ (orange circles) and the linear interpolation $x_{s}=\underline{x}_{s+1}$ (green squares) obtained from a coarser-scale discretization. The top figure shows Eq.˜7, where $x_{s+1}=x^{*}_{s+1}$ is exact. The bottom figure shows Eq.˜8, where $x_{s+1}=x^{*}_{s+1}+\delta$ contains error.
4.2 Greedy Multiscale
We use the following approach to show convergence of the greedy multiscale algorithm described in Algorithm˜2. At each scale $s$, we use the iterate convergence (Definition˜2.6) to bound the error between our iterate $x_{s}^{K_{s}}$ and the solution $x_{s}^{*}$ at that scale. Then we use the inexact interpolation lemma (Lemma˜4.7) to bound the error between our iterate $x_{s}^{K_{s}}$ and interpolation $\underline{x}_{S}^{K_{S}}$. This interpolated iterate becomes the initialization $x_{S-1}^{0}=\underline{x}_{S}^{K_{S}}$ at the next finest scale.
Theorem 4.9 (Greedy Multiscale Error Bound).
Let $q$ denote the rate of convergence of an algorithm applied to a loss function $\tilde{\mathcal{L}}$, let $L_{f^{*}}$ be the Lipschitz constant of the solution function $f^{*}$, and let $S$ denote the coarsest scale. The greedy multiscale method in Algorithm˜2 returns iterate $x_{1}^{K_{1}}$ at the finest scale satisfying
$\left\lVert x^{K_{1}}_{1}-x^{*}_{1}\right\rVert_{2}\leq\sqrt{2^{S-1}}(q)^{r_{S}}\left\lVert x_{S}^{0}-x_{S}^{*}\right\rVert_{2}+L_{f^{*}}\frac{\left\lvert u-\ell\right\rvert}{2\sqrt{2^{S+1}}}\sum_{s=1}^{S-1}2^{s}(q)^{r_{s}},$
where $r_{s}=\sum_{t=1}^{s}K_{t}$ denotes the cumulative iteration count through scale $s$.
Proof 4.10.
See LABEL:sec-greedy-multiscale-descent-error-proof.
4.3 Lazy Multiscale
We establish convergence of the lazy multiscale algorithm using the framework from Theorem˜4.9. The key modification is to freeze the interpolated values $\underline{x}_{s}^{K_{s}}$ that correspond to points from the previous iteration $x_{s}^{K_{s}}$. The algorithm updates only the newly interpolated values, denoted $x_{(s)}^{k}$ (see Definition˜2.3).
Theorem 4.11 (Lazy Multiscale Error Bound).
Let $q$ denote the convergence rate of an algorithm with iterate convergence on a loss function $\tilde{\mathcal{L}}$. Under the setting of Section˜4.1, the lazy multiscale method satisfies
$\left\lVert x_{1}^{K_{1}}-x_{1}^{*}\right\rVert$ $\leq\left\lVert x_{S}^{0}-x_{S}^{*}\right\rVert(q)^{K_{S}}\prod_{s=1}^{S-1}(1+(q)^{K_{s}})$
$\qquad+\ \tfrac{1}{2}L_{f^{*}}\left\lvert u-\ell\right\rvert\sum_{s=1}^{S-1}\frac{1}{\sqrt{2^{S-s}}}(q)^{K_{s}}\prod_{j=1}^{s-1}\left(1+(q)^{K_{j}}\right).$
When using constant iterations $K_{s}=K$ at each scale, this reduces to
$\left\lVert x_{1}^{K_{1}}-x_{1}^{*}\right\rVert$ $\leq\left\lVert x_{S}^{0}-x_{S}^{*}\right\rVert d(K)(d(K)+1)^{S-1}$
$\qquad+\ \tfrac{1}{2}L_{f^{*}}\left\lvert u-\ell\right\rvert\frac{d(K)\left(\sqrt{2^{S}}\left(d(K)+1\right)^{S}-\sqrt{2}\left(d(K)+1\right)\right)}{\sqrt{2^{S}}\left(d(K)+1\right)\left(\sqrt{2}d(K)+\sqrt{2}-1\right)}$
where $d(K)=(q)^{K}\in(0,1)$.
Proof 4.12.
See LABEL:sec-lazy-multiscale-descent-error-proof.
We summarize the convergence with Corollary˜4.13.
Corollary 4.13 (Multiscale Convergence).
[…]
5 Relation Between the Discretized and Continuous Problems
[…]
We assume $t\in\mathbb{R}^{I}$ forms a uniform grid on the interval $\ell,u]$ with spacing $\Delta t=t[{i+1}]-t[i]=\frac{u-\ell}{I-1}$.
Lemma 5.1 (Distance Between Piecewise Linear Functions).
Let $f,g:\mathbb{R}\to\mathbb{R}$, and let $x,y\in\mathbb{R}^{I}$ be samples $x[i]=f(t[i])$ and $y[i]=g(t[i])$. Construct piecewise linear approximations (Definition˜2.4) $\hat{f}$ of $f$ and $\hat{g}$ of $g$. Then
$\|\hat{f}-\hat{g}\|_{2}^{2}\leq\tfrac{2}{3}\Delta t\|x-y\|_{2}^{2}.$
Proof 5.2.
See LABEL:sec-piecewise-function-distance-proof.
We bound the $L^{2}$ distance between a Lipschitz function and its piecewise linear approximation.
Lemma 5.3 (Piecewise Linear Function Approximation).
Let $f$ be an $L_{f}$-Lipschitz function on $\ell,u]$, and let $\hat{f}$ denote its piecewise linear approximation (Definition˜2.4) on the uniform grid $t\in\mathbb{R}^{I}$. Then
$\|\hat{f}-f\|_{2}^{2}\leq\tfrac{2}{15}(u-\ell)\Delta t^{2}L_{f}^{2}.$
Proof 5.4.
See LABEL:sec-piecewise-function-approximation-proof.
Lemma˜5.3 compares favourably with existing results. De Boor [de_boor_practical_2001, Ch. III, Eq. 17] establishes the bound
$\|\hat{f}-f\|_{2}^{2}\leq\tfrac{1}{16}\Delta t^{4}\|f^{\prime\prime}\|_{2}^{2},$
which converges faster as the interval width $\Delta t\to 0$ (equivalently, $I\to\infty$), but requires $f\in C^{2}([\ell,u])$ with bounded second derivative. Lipschitz continuity and second-order differentiability doesn’t suffice: for example, the parametrized soft-plus function $f(x)=\ln(1+\exp(cx))/c$ is $1$-Lipschitz for all $c>0$, but has unbounded second derivative $f^{\prime\prime}(0)=c/2$ as $c\to\infty$.
De Boor showed that the error vanishes for continuous functions (not necessarily Lipschitz), though possibly a slower rate than Lemma˜5.3 [de_boor_practical_2001, Ch. III, Eq. 18]. Kunoth et al. [kunoth_splines_2018, Sec. 1.5, Th. 18] generalizes the approximation error to functions in Sobolev spaces, achieving for differentiable Lipschitz functions a rate comparable to Lemma˜5.3:
(9) $\|\hat{f}-f\|_{2}^{2}\leq K\Delta t^{2}\|f^{\prime}\|_{2}^{2}$}
for some constant $K>0$.
We cannot expect a better rate because Lipschitz functions are differentiable almost everywhere [heinonen_lectures_2004]. For Lipschitz functions (possibly not differentiable), the derivative, where it exists, satisfies $\left\lvert f^{\prime}(x)\right\rvert\leq L_{f}$, giving $\left\lVert f^{\prime}\right\rVert_{2}^{2}\leq(u-\ell)L_{f}^{2}$. Thus, Lemma˜5.3 explicitly calculates the constant $K=2/15$ in Eq.˜9.
Theorem 5.5 (Connection Between the Continuous and Discrete Problems).
Let $f:[\ell,u]\to\mathbb{R}$ be $L_{f}$-Lipschitz with discretization $x^{*}\in\mathbb{R}^{I}$. Let $x\in\mathbb{R}^{I}$, and construct the corresponding piecewise linear function $\hat{f}_{x}$ (Definition˜2.4). For $\epsilon>0$, if
$I>C\epsilon^{-1}+1\quad\text{and}\quad\|x-x^{*}\|_{2}^{2}<D\epsilon$
then
$\|\hat{f}_{x}-f\|_{2}<\epsilon,$
where $C=\sqrt{\frac{8}{15}}(u-\ell)^{3/2}L_{f}$ and $D=\sqrt{\frac{3}{40}}(u-\ell)^{1/2}L_{f}$.
Proof 5.6.
See LABEL:sec-problem-connection-proof.
[…]
6 Comparison with Projected Gradient Descent
[…]
Lemma˜2.7 establishes iterate convergence (cf. Definition˜2.6) for projected gradient descent. We use this to derive expected convergence bounds for three approaches: projected gradient descent at scale $s=1$ (Theorem˜6.1), greedy multiscale descent (Theorem˜6.4), and lazy multiscale descent (Theorem˜6.14). After presenting these results, we demonstrate that both multiscale variants achieve tighter error bounds with lower computational cost when the problem size and Lipschitz constant are sufficiently large.
Theorem 6.1 (Expected Projected Gradient Descent Convergence).
[…]
(10) $\mathbb{E}\left\lVert x_{1}^{K}-x_{1}^{*}\right\rVert\leq(q)^{K}\sqrt{2^{S}+2}.$}
Corollary 6.2 (Expected Number of Iterations).
If we iterate projected gradient descent $K$ times, where
$K\geq\frac{\log(1/\epsilon)+{\log(2^{S}+2)}/{2}}{-\log q},$
then the expected error is bounded as
$\mathbb{E}\left\lVert x_{1}^{k}-x_{1}^{*}\right\rVert_{2}\leq\epsilon.$
Proof 6.3.
See LABEL:sec-expected-pgd-convergence-proof.
Theorem 6.4 (Expected Greedy Multiscale Convergence).
Assume the same setting as Theorem˜6.1, but use the greedy multiscale descent method starting at scale $s=S$. At the coarsest scale, the expected initial error satisfies
$\mathbb{E}\left\lVert x_{S}^{0}-x_{S}^{*}\right\rVert_{2}\leq 2.$
Performing $K_{s}$ iterations at each scale yields the expected final error
(11) $\mathbb{E}\,\|x_{1}^{K_{1}}-x_{1}^{*}\|_{2}\leq(q)^{K_{1}}\sqrt{2^{S+1}}\left((q)^{r_{S}}+\frac{L_{f^{*}}\left\lvert u-\ell\right\rvert}{{2^{S+2}}}\sum_{s=1}^{S-1}2^{s}(q)^{r_{s}}\right)$}
where $r_{s}=\sum_{t=2}^{s}K_{t}$ with the base case $r_{1}=0$.
Proof 6.5.
See LABEL:sec-expected-greedy-convergence-proof.
Determining the iteration count required to achieve a specified accuracy is less direct for multiscale methods than for projected gradient descent. Moreover, a fair comparison requires accounting for computational cost: an iteration at coarse scale $S$ requires fewer floating-point operations than an iteration at the finest scale $s=1$. We therefore configure each iteration of multiscale to simultaneously achieve lower total cost than fine-scale-only optimization while obtaining a tighter expected error bound. Specifically, we require Eq.˜11 to be smaller than Eq.˜10. This requires costing multiscale in terms of one projected gradient descent iteration at the finest scale $s=1$.
Lemma 6.6 (Cost of Greedy Multiscale).
Suppose projected gradient descent with $K$ iterations at the finest scale $s=1$ has total cost $C_{\mathrm{GD}}=AK$, where $A>0$ is the cost per iteration at that scale. Assume the cost of one projected gradient descent iteration scales polynomially with problem size as $C_{s}=C(I_{s})^{p}$ for $I_{s}=2^{S-s+1}+1$ points and constants $C>0$ and $p\geq 1$. Then greedy multiscale descent with $K_{s}$ iterations at scale $s$ has total cost $C_{\mathrm{GM}}$ that satisfies
$C_{\mathrm{GM}}\leq A\sum_{s=1}^{S}K_{s}\left(\frac{3}{5}\right)^{s-1}.$
Proof 6.7.
See LABEL:sec-cost-of-gd-vs-ms-proof.
The factor $3/5$ is the ratio of the number of points at the two coarsest scales: $I_{S}/I_{S-1}=3/5$.
The assumption that iteration cost scales as $CI^{p}$ for $C>0$ and $p\geq 1$ holds for common projects and gradient computations. For example, projection of a vector $y\in\mathbb{R}^{I}$ onto the sphere $\mathcal{S}^{I-1}$, given by
[…]
scales linearly in II ($p=1$). The least-squares gradient update for $A\in\mathbb{R}^{I\times I}$,
[…]
scales quadratically in II ($p=2$).
Lemma˜6.6 establishes a selection of iterations counts $K_{s}$ at each scale such that greedy multiscale is cheaper than projected gradient descent.
Corollary 6.8 (Conditions for Greedy Multiscale to be Cheaper).
Consider projected gradient descent at scale $s=1$ with $K$ iterations. Greedy multiscale with $K_{s}$ iterations at scale $s$ achieves lower cost under either iteration plan:
  • $K_{s}=\left\lceil 2K/5\right\rceil-1$ iterations at every scale; or
  • $K_{s}=1$ iterations at every scale except the finest scale, where $K_{1}=K-2$.
Proof 6.9.
This follows from Lemma˜6.6 combined with bounds on the sums
$\sum_{s=1}^{S}\left(\frac{3}{5}\right)^{s-1}=\frac{1-(3/5)^{S}}{1-(3/5)}<5/2$
and
$\sum_{s=2}^{S}\left(\frac{3}{5}\right)^{s-1}=\frac{3/5-(3/5)^{S}}{1-(3/5)}<3/2<2.$
The analysis in Corollary˜6.8 ignores the cost of interpolation and memory allocation when changing iterate dimensions between scales. These costs become negligible as the iteration counts increase. Empirical results in Section˜7 confirm that greedy multiscale remains cheaper in practice.
Having established conditions that ensure greedy multiscale is cheaper than projected gradient descent, we now establish conditions that ensure tighter expected error bounds.
Corollary 6.10 (Sufficient Conditions for Greedy Multiscale to be Better).
Assume the base algorithm has iterate convergence (Definition˜2.6) with rate $q\in(0,1/2)$. Assume the finest scale problem has $I=2^{S}+1$ points, where the number of scales $S$ satisfies
$S\geq\max\left\{4,\,\log_{2}\left(\frac{L_{f^{*}}\left\lvert u-\ell\right\rvert}{\sqrt{2}(q)^{2}(1-2q)(1-\sqrt{2}q)}\right)\right\}.$
Then greedy multiscale with one iteration $K_{s}=1$ at each scale except the finest (where $K_{1}=K-2$) simultaneously achieves lower total cost and a tighter expected final error bound than projected gradient descent with $K$ iterations.
Proof 6.11.
See LABEL:sec-greedy-cost-proof.
We establish analogous results for lazy multiscale.
Lemma 6.12 (Cost of lazy Multiscale).
Suppose projected gradient descent with $K$ iterations has total cost $C_{\mathrm{GD}}=AK$, where $A>0$ is the cost per iteration at the finest scale $s=1$. Under the polynomial cost scaling assumption $C_{s}=\Theta(I_{s}^{p})$ for $I_{s}=2^{S-s+1}+1$ points and $p\geq 1$, lazy multiscale descent with $K_{s}$ iterations at scale $s$ has total cost $C_{\mathrm{LM}}$ that satisfies
$C_{\mathrm{LM}}\leq A\left(\sum_{s=1}^{S-1}2^{-s}K_{s}+2^{-S}3K_{S}\right).$
With $\left\lceil\frac{4}{5}K\right\rceil-1$ iterations at each scale,
$C_{\mathrm{LM}}<C_{\mathrm{GD}}.$
Proof 6.13.
See LABEL:sec-cost-of-gd-vs-lazy-ms-proof
Theorem 6.14 (Expected Lazy Multiscale Convergence).
Assume the same setting as Theorem˜6.1, but use the lazy multiscale starting at coarsest scale $S$ with $I=2^{S}+1$ points at the finest scale. The expected initial error satisfies
$\mathbb{E}\left\lVert x_{S}^{0}-x_{S}^{*}\right\rVert_{2}=2.$
Performing constant $K_{s}=K$ iterations at each scale yields the expected final error
$\mathbb{E}\left\lVert x_{1}^{K}-x_{1}^{*}\right\rVert_{2}$
$\leq d(K)\left(2(d(K)+1)^{S-1}+\frac{L_{f^{*}}}{2}\left\lvert u-\ell\right\rvert\frac{\left(\sqrt{2^{S}}\left(d(K)+1\right)^{S}-\sqrt{2}\left(d(K)+1\right)\right)}{\sqrt{2^{S}}\left(d(K)+1\right)\left(\sqrt{2}d(K)+\sqrt{2}-1\right)}\right),$
where $d(K)=(q)^{K}$ and the solution function $f^{*}$ is $L_{f^{*}}$-Lipschitz on $\ell,u]$.
Proof 6.15.
See LABEL:sec-expected-lazy-convergence-proof.
We establish sufficient conditions that ensure lazy multiscale achieves both lower cost and tighter error bounds than projected gradient descent.
Corollary 6.16 (Sufficient Conditions for Lazy Multiscale to be Better).
Assume the base algorithm has iterate convergence (Definition˜2.6) with rate $q\in(0,1)$. Assume projected gradient decent runs for $K>5(\log_{q}(\sqrt{2}-1)-1)/4$ iterations, and lazy multiscale runs for $K^{\prime}=\lceil 4K/5\rceil-1$ iterations at every scale. If the finest scale has at least $I=2^{S}+1$ points where
$S\geq\frac{\log\left(\frac{q^{-K/5-1}}{1+q^{K^{\prime}}}\left(2+\frac{L_{f^{*}}\left\lvert u-\ell\right\rvert}{2(\sqrt{2}(1+q^{K^{\prime}})-1)}\right)\right)}{\log\left(\frac{\sqrt{2}}{1+q^{K^{\prime}}}\right)},$
then lazy multiscale is cheaper and achieves a tighter expected error bounds than projected gradient descent.
Proof 6.17.
See LABEL:sec-lazy-cost-proof.
Corollaries˜6.10 and 6.16 reveal several insights. Both multiscale variants require a minimum number of scales $S$ (equivalently a minimum sized problem $I$) for the multiscale optimization to outperform single scale methods. Theorem˜5.5 establishes that arbitrarily small error between continuous and discretized solutions requires arbitrarily large problem size $I$. Both corollaries show that the required $I$ increases with the Lipschitz constant $L_{f^{*}}$ and the domain size $\left\lvert u-\ell\right\rvert$, consistent with Theorem˜5.5. A subtle requirement is that the base algorithm cannot have iterative convergence with $q=0$, which would enable solving the discretized problem in one iteration, and thus eliminate the benefit of multiscale.
Corollaries˜6.10 and 6.16 differ in their constraints on the convergence rate $q$. Greedy multiscale with one iteration per scale requires convergence rate $q\in(0,1/2)$ because interpolation error must remain smaller than the algorithmic progress at each scale. Lazy multiscale, on the other hand, allows $q\in(0,1)$ but requires a minimum of $K^{\prime}>\log_{q}(\sqrt{2}-1)$ iterations at each scale to sufficiently control error accumulation. This is a mild requirement: two iterations per scale suffice for lazy multiscale to improve on the base algorithm when $0<q<(\sqrt{2}-1)^{1/2}\approx 0.6436$, and both algorithms require arbitrarily many iterations to drive the final error to zero.
7 Numerical Experiments and Benchmarks
We evaluate the greedy multiscale method on two density demixing problems and compare it to standard projected gradient descent. These problems seek to recover mixtures of continuous probability densities from noisy samples. We consider synthetic data in Section˜7.2 and geological survey data in Section˜7.3.
The project-gradient multiscale and single-scale algorithms are implemented in the Julia package BlockTensorFactorization.jl [Richardson_BlockTensorFactorization_jl]: the function factorize implements projected gradient descent for the Tucker-1 decomposition problem at a single scale, and multiscale_factorize implements the greedy multiscale method (Algorithm˜2). All experiments ran on an Intel Core i7-1185G7 with 32GB of RAM, without parallelization, matching the hardware used for the example experiment in Section˜2.3.
7.1 Density Demixing as Optimization
Graham et. al. [graham_tracing_2025] provides a full treatment of this formulation which we summarize here. Additional details are given in LABEL:additional-tensor-details-and-reference.
Given $I$ mixtures $y_{i}:\mathcal{D}\subseteq\mathbb{R}^{N}\to\mathbb{R}_{+}$ of $R$ source probability density functions $b_{r}:\mathcal{D}\subseteq\mathbb{R}^{N}\to\mathbb{R}_{+}$,
$y_{i}(x)=\textstyle\sum_{r\in[R]}a_{i,r}b_{r}(x),$
we seek to recover the sources $\{b_{r}\}$ and their mixing coefficients $a_{i,r}$. We assume $R<I$ and that the source densities $\{b_{r}\}$ are linearly independent, which ensures this is a well-posed problem.
The continuous problem formulation is
$\min_{\{a_{i,r}\},\{b_{r}\}}\tfrac{1}{2}\textstyle\sum_{i\in[I]}\left\lVert\sum_{r\in[R]}a_{i,r}b_{r}-y_{i}\right\rVert_{2}^{2}$
such that for all $i\in[I]$,
(12a) $1$ $=\textstyle\sum_{r\in[R]}a_{i,r},\qquad$ $a_{i,r}$ $\geq 0,$
(12b) $1$ $=\textstyle\int_{\mathcal{D}}b_{r}(x)\,dx,\qquad$ $b_{r}(x)$ $\geq 0\quad\forall x\in\mathcal{D},\ \forall r\in[R].$
Discretizing yields the Tucker-1 decomposition problem (see LABEL:def-tucker-1-decomposition)
[…]
where
[…]
In the geological data example (Section˜7.3), each dimension of $b_{r}$ is independent. This allows a simpler discretization, so that $B$ becomes a third-order tensor with
[…]
and we reinterpret $Y\in\mathbb{R}_{+}^{I\times J\times K}$ accordingly.
7.2 Synthetic Data
We generate three source distributions for a synthetic test. Each distribution is a 3-dimensional product of standard continuous distributions; see LABEL:sec-synthetic-data-setup for details, with full code available in this paper’s GitHub repository [Richardson_multiscale_paper_code]. Benchmarking the two approaches yields the following for the single-scale factorize function:
B
Range (min...max): 1.305 s...3.487 s   GC (min...max): 15.07%
Time (median): 2.412 s   GC (median): 27.67%
Time (mean $\scriptstyle\pm$} $\sigma$): 2.418 s $\scriptstyle\pm$} 628.420 ms   GC (mean $\scriptstyle\pm$} $\sigma$): 25.08%
Memory estimate: 1.46 GiB, allocs estimate: 2189884.
[…]
and for the factorization at multiple scales with the multiscale_factorize function:
B
Range (min...max): 231.299 ms...1.583 s   GC (min...max): 8.10%
Time (median): 334.698 ms   GC (median): 11.61%
Time (mean $\scriptstyle\pm$} $\sigma$): 455.869 ms $\scriptstyle\pm$} 313.358 ms   GC (mean $\scriptstyle\pm$} $\sigma$): 23.90%
Memory estimate: 366.81 MiB, allocs estimate: 569530.
[…]
The function multiscale_factorize runs roughly seven times faster and uses about one-fourth of the memory. This comparison includes overhead such as interpolating the tensor and repeatedly calling the internal factorize function, which represents approximately $4\%$ of the total time. The specific numbers are less important than the qualitative comparison: both algorithms could be optimized and benchmarked on state-of-the-art hardware, but this basic implementation demonstrates that multiscale methods can accelerate single-scaled algorithm.
In these multiscale_factorize benchmarks, rather than performing one iteration at each scale coarser than the finest scale, we advance to the next scale once sufficient progress has been made; see LABEL:sec-synthetic-data-setup for details. This differs from our analysis in Corollary˜6.10, which suggests using exactly $K_{s}=1$ at coarse scales. However, running additional iterations at coarser scales often reduces the total number of iterations required at the finest scale.
Figure˜8 (left) shows median runtime as a function of the number of fixed coarse iterations $K_{s}$ in multiscale for $s=S,S-1,\dots,2$, where we run iterations at the finest scale until the objective falls below $10^{-6}$. This figure demonstrates that performing more than one iteration at coarser scales improves multiscale performance, but only to a point: in Fig.˜8 (left), more than 12 iterations at each coarse scale wastes time. Figure˜8 (right) illustrates this by showing the number of fine-scale iterations $K_{1}$ needed to converge to a minimum value. Our implementation of multiscale_factorize [Richardson_BlockTensorFactorization_jl] addresses this by allowing multiple iterations at coarser scales and advancing to the next scale when other criteria are met, such as small gradients or objective values.
Both plots in Fig.˜8 show that multiscale with exactly $K_{s}=1$ iterations at coarse scales $s=S,S-1,\dots,2$ performs slightly worse than single-scale optimization. We attribute this to interpolation and to overhead costs in the multiscale approach. The problem size of $2^{6}+1=65$ points along each continuous dimension may be too small to see immediate benefit from multiscale when only one iteration is performed at each scale; larger problems may needed to realize the advantage illustrated by Fig.˜1. However, flexible criteria to advance between scales may make multiscale competitive even for smaller problem.
Figure 8: (Left) Median time until convergence at the finest scale $s=1$ as a function of number of fixed iterations $K_{s}$ at coarser scales $s=S,S-1,\dots,2$. (Right) Median number of iterations $K_{1}$ at the finest scale $s=1$. Shaded ribbon shows the bottom and top quartile times (left) and iterations (right).
7.3 Real Geological Data
We use the same sedimentary data and Tucker-$1$ model described in Graham et. al. [graham_tracing_2025] to factorize a tensor containing mixtures of estimated densities. LABEL:sec-geo-numerics-details provides the details, and the full code appears in distribution_unmixing_geology.jl in this paper’s GitHub repository [Richardson_multiscale_paper_code]. For factorize, we obtain
B
Range (min...max): 174.316 ms...593.727 ms   GC (min...max): 19.75%
Time (median): 416.816 ms   GC (median): 19.91%
Time (mean $\scriptstyle\pm$} $\sigma$): 389.401 ms $\scriptstyle\pm$} 108.955 ms   GC (mean $\scriptstyle\pm$} $\sigma$): 19.90%
Memory estimate: 361.43 MiB, allocs estimate: 139855.
[…]
and for multiscale_factorize, we observe
B
Range (min...max): 78.320 ms...153.810 ms   GC (min...max): 0.00%
Time (median): 91.376 ms   GC (median): 0.00%
Time (mean $\scriptstyle\pm$} $\sigma$): 95.542 ms $\scriptstyle\pm$} 19.154 ms   GC (mean $\scriptstyle\pm$} $\sigma$): 11.24%
Memory estimate: 105.80 MiB, allocs estimate: 1444133.
[…]
By every metric, the multiscale approach is faster and uses less memory than the standard single-scale factorization. Comparing median time and memory estimates, the multiscale method runs roughly four times as fast and uses about a third of the memory. These wall clock-times include significant garbage collection (GC), suggesting that design improvements such as preallocated arrays could reduce memory usage further. Such improvements would accelerate both factorize and multiscale_factorize. Nevertheless, we observe that multiscale_factorize triggers garbage collection less frequently because more computations occur on smaller arrays.
8 Conclusion
We developed a multiscale approach for optimizing Lipschitz continuous functions by solving discretized problems at progressively finer scales. Our convergence analysis establishes that both greedy and lazy variants achieve the finest-scale solution through per-scale descent combined with controlled interpolation error. The framework extends to linear and norm constraints and to tensor-valued functions.
The explicit error bounds that we derive quantify the tradeoff between discretization granularity and solution accuracy, and prove that multiscale can outperform direct fine-scale optimization when the problem size and Lipschitz constant are sufficiently large.
Several directions merit further investigation. Our analysis assumes a fixed number of base algorithm iterations at each scale. However, an adaptive iteration count that responds to convergence progress could improve efficiency. With regard to the discretization approach, the dyadic scheme might be replaced by adaptive grid refinement based on local Lipschitz estimates. Second-order base algorithms present another avenue, though their cost would be better characterized in terms of linear system solves rather than gradient evaluations. Finally, interpreting the pointwise discretization as a discrete wavelet transform suggests generalizations to other wavelet expansions.